OpenStudy (mikey):

Bob and Susie wash cars for extra money over the summer. Bob's income is determined by f(x) = 6x + 13, where x is the number of hours. Susie's income is g(x) = 4x + 18. If Bob and Susie were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Bob works 3 hours. Create the function h(x), and indicate if Bob will make more money working alone or by teaming with Susie. h(x) = 2x + 5, work alone h(x) = 2x + 5, team with Susie h(x) = 10x + 31, team with Susie h(x) = 10x + 31, work alone

OpenStudy (francesca_marie):

h(x) = 2x + 5, work alone h(x) = 2x + 5, team with Susie h(x) = 10x + 31, team with Susie h(x) = 10x + 31, work alone

OpenStudy (mikey):

@3mar

OpenStudy (francesca_marie):

._. kay den

OpenStudy (mikey):

what u mean kay den @Francesca_Marie

OpenStudy (mikey):

@Awolflover1

OpenStudy (mikey):

@AloneS

OpenStudy (francesca_marie):

I helped you and you tagged someone else xD

OpenStudy (mikey):

u just rewrote the choices and I was tagging someone that I knew helped me before, sorry

OpenStudy (francesca_marie):

oh i meant to put h(x) = 10x + 31, team with Susie was correct xD

OpenStudy (3mar):

Really I am confused about it. Can we think together?

OpenStudy (mikey):

I was extremely confused as well

OpenStudy (francesca_marie):

Yeah

OpenStudy (francesca_marie):

well x = 3 and f(x) is Bob's income function

OpenStudy (francesca_marie):

so f(3) represents how much money he earns alone

OpenStudy (3mar):

great f(3)=31 then?

OpenStudy (francesca_marie):

sorry my computer turned off lol

OpenStudy (3mar):

combined: h(x)=f(x)+g(x) =6x+13+4x+18 =10x+31 for h(3)=30+31=61 So with team is better the third choice as Francesca_Marie said. He was right!

OpenStudy (francesca_marie):

She* xD

OpenStudy (3mar):

Sorry, names do not tell me who are they!! Nice work, Maria